# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v2

@@ Abstract
The technological development of increasingly larger quantum processors on different quantum platforms raises the problem of how to fairly compare their performance, known as quantum benchmarking of quantum processors. This is a challenge that computer scientists have already faced when comparing classical processors, leading to the development of various mathematical tools to address it, but also to the identification of the limits of this problem. In this work, we briefly review the most important aspects of both classical processor benchmarks and the metrics comprising them, providing precise definitions and analyzing the quality attributes that they should exhibit. Subsequently, we analyze the intrinsic properties that characterize the paradigm of quantum computing and hinder the naive transfer of strategies from classical benchmarking. However, we can still leverage some of the lessons learned such as the quality attributes of a good benchmark. Additionally, we review some of the most important metrics and benchmarks for quantum processors proposed in the literature, assessing what quality attributes they fulfill. Finally, we propose general guidelines for quantum benchmarking. These guidelines aim to pave the way for establishing a roadmap towards standardizing the performance evaluation of quantum devices, ultimately leading to the creation of an organization akin to the Standard Performance Evaluation Corporation (SPEC).
@@ I Introduction
The lack of standardized benchmarking rules, together with the increasing competition among major industry players vying to outperform their rivals in the quantum computing race, might lead to a situation that resonates with Goodhart’s law goodhart1975problems : ”Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes”. In other words, the absence of a rigorously formulated and scrutinized approach to benchmarking in quantum computing may distort research priorities and hinder the development of truly scalable and reliable quantum processors. This bias occurs due to benchmarking models playing a crucial role on favouring what portion of the total ”behavior space” is captured SkadronChallengesInComputerArchitectureEvaluation .
@@ II What can we learn from Classical benchmarking?
A natural question that then arises is what we mean by performance. Performance is typically conceived in the classical computing scope as the useful work that a computer can get done in a given amount of time, and/or with a set of given resources. Indeed, in the past, it was common to hold the premise that ”the time required to perform a specified amount of computation is the ultimate measure of computer performance” CharacterizingComputerPerformanceWithASingleDigit . For classical computers, there have been historically two major approaches to the performance analysis problem: (1) Experimenting and measuring the actual system, and (2) Experimenting with a model of the system. For the latter, for any kind of performance evaluation not limited to computing devices you can use a physical model or a mathematical model. For computational performance assessment, we prefer to build models that can represent the system in terms of quantitative and logical relations. These models can later be either solved using mathematical tooling to obtain an analytical solution, or simulated Zeigler2018fg ; SimulationModelingAndAnalysisLawAverill . Even though in this work we will focus on the measurements of experiments with the actual systems, it is possible that we will see other approaches involving system modeling in the future. So far, and to the best of our knowledge, models have only been used by the quantum community for the assessment of quantum networks Wu2021 ; Coopmans2021 , still there are other event-based models of quantum phenomena published DeRaedt2005 ; Willsch2020 .
@@ II.1 Characteristics of a good benchmark
To be able to build a good benchmark, we introduce below a list of quality traits to ensure that it is comprehensive, accurate, and representative of system performance. These attributes are in line with those proposed by Kounev Kounev2020mh and also reflect the most desirable characteristics of benchmarks proposed by other authors in the field HowToBuildABenchmark ; GustafsonAndSnell_HINT ; Henning_SpecCpu2000 ; Huppler_TheArtOfBuildingGoodBenchmark ; Sim_BenchmarkingToAdvanceResearch .
@@
Similarly, verifiability presents comparable challenges that we should carefully consider too. In industry, benchmarks are typically run by vendors who have a vested interest in the outcomes. In contracts, in academia results are subjected to a peer review process to be repeated and built upon by other researchers. Good benchmarks should introduce self-validation, since this may improve verifiability and usability, ensuring that results can be independently corroborated and effectively utilized.
@@
These statements highlight the susceptibility of benchmarks to manipulation, emphasizing the need for cautious interpretation and critical evaluation when drawing conclusions based on the results from ratio games. Other researchers such as Mashey have also call out the difficulties of aggregating data, and the controversy of using one one Pythagorean mean over the others, as well as stated the common agreement that ”performance is not a single number” Mashey2004 . Nonetheless, even though there is agreement on the fact that a single number is delusional, there are still discussions to determine what is the preferred metric, sparking debates over which numerical representations are deemed more optimal in each context.
Finally, it is important to remember that while a good benchmark can provide a reliable and accurate measure of system performance, it cannot solve the problem of evaluating conflicting aspects of a computer device. Instead, it can help to properly balance the trade-offs for a particular objective and influence its strengths and weaknesses. It is not possible for a single benchmark to excel in all areas, and there is no silver bullet to solve this problem SkadronChallengesInComputerArchitectureEvaluation . Therefore, multiple workloads and benchmarks are necessary to comprehensively evaluate the performance of a system. Each benchmark may focus on different aspects of performance and may use different metrics and methodologies. At best, we can obtain a set of benchmarks that will be representative of the behavior space we want to model and measure. By combining the results of these multiple benchmarks, it would be possible to obtain a comprehensive and accurate picture of the system. Ultimately, the choice of benchmark(s) should depend on the specific use case and the objectives of the evaluation, as the relevancy requirement already stated.
@@ II.2 What types of benchmarks are there?
Benchmarks can be classified according to multiple criterias, ranging from the objective of what portion of the behavior space they want to assess to the way the workload is defined. All of these categories fall under one of the following three different strategies when trying to assess the performance Lilja2005o ; Kounev2020mh :
@@ II.3 What makes a good metric
    Reliable. It must rank systems consistently with respect to a property that is subject to evaluation. In other words, if system $A$ performs better than system $B$ with respect to the property under evaluation, then the values of the metric for the two systems should consistently indicate so. The reliability of a metric is closely correlated with the relevancy of a benchmark.
@@
    Consistent. It has the same units and the same precise definition across different systems or configurations of the same system. When this property is not met, the benchmark will suffer poor verifiability and potentially lack of relevancy.
@@
Finally, it is important to notice that the simple act of measuring a property using a metric might be tough. The effect of the Simpsons’ paradox serves as an adage to the designers of benchmarks and metrics, calling out the possibility of reversing or disappearing the trend on the data when combining it leboudec2010performance .
@@ II.4 Examples of classical benchmarks
Other examples of controversial benchmarks used historically are the synthetic ones, such as the Whetstone WhetstoneBook . Even though it was proposed to measure the floating point performance of a CPU, it can be misleading when compared with real-life situations. Nowadays, computers can easily fit the main loop of Whetstone in the low-level caches of the CPU, keeping the math coprocessor pipeline filled and achieving its optimal speed. Moreover, they face the issue of most synthetic benchmarks, they are not guaranteed to be correlated with real application scenarios, and fail to meet true relevancy.
@@ III Quantum benchmarking
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion QCion ; Zhang_2017 , superconducting circuits Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 , photonic QCPhoton ; PhysRevLett.127.180502 , and cold atoms Bernien2017ProbingMD ; Ebadi2020QuantumPO . Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including quantum annealing for optimization problems Rajak_2022 ; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems Kendon_2010 ; and digital quantum computing, which uses a gate-based quantum devices to simulate a quantum system doi:10.1126/science.273.5278.1073 . These three paradigms offer different perspectives on the practical applications of quantum computing.
@@
In the quantum computing field, we have deceptively misleadingly used the term benchmarking. While it traditionally refers to a series of tests evaluating the performance of a device, its application has expanded to encompass the metrics evaluated as part of the test, the actual tests, or even for verifying the success of a test. We believe it is crucial to establish a unified terminology for quantum benchmarking to provide a solid foundation for comprehensive and effective evaluations of quantum systems.
@@ III.1 Nuances of benchmarking quantum computers
    Fault Tolerant Quantum Computers. Once sufficiently large-distance codes may be implemented, then there will be theoretically no limit in the size of the quantum circuits and the full computational power of quantum computers will be deployed.
@@
    Fully Functional Quantum Computers. Existing quantum algorithms and circuits consider the use of ancillary qubits, which play a similar role as cache memory in classical processors. However, for Fully Functional Quantum Computers, there will also be necessary the existence of Quantum Random Access Memories (qRAM) QuantumRam , and even long-term (error-corrected) quantum memories Terhal2015 .
@@
The challenges that we are facing today within the NISQ era are different from the ones we will be facing in the upcoming eras, and hence the benchmarks should reflect this. Given the wide diversity of technological platforms and paradigms for implementing quantum computers, it is likely that benchmarks in the NISQ era will focus on comparing the performance among different quantum computers within subsets of the broad existing variety. As technology advances, it is expected that the wide variety of existing technological platforms will naturally decrease, preserving only those that are most competitive for the implementation of quantum computing. In this context, as technology continues to progress, comparing the performance of quantum computers implemented on different technological platforms will become more feasible. On the other hand, achieving a fair comparison of performance between quantum computers and classical computers at present is not viable, as quantum computers require different workloads, metrics, and overall approaches to benchmarking than classical computers. This is expected to occur in the era of Fully Functional Quantum Computers. In Sec. IV, we will discuss further what recommendations we suggest to properly face the different quantum technological development eras.
@@
  • The fundamental unit of information in a quantum computer is the qubit (quantum bit). While classical bits can only represent a single binary value 0 or 1, qubits have the ability to exist in multiple states simultaneously, thanks to a phenomenon known as superposition. As a result, a qubit can be expressed as a linear combination of both the 0 and 1 states.
@@
Apart from these are fundamental features associated to quantum physics, there are other relevant aspects of quantum computing which mainly affect the performance for the NISQ and PQEC eras.
@@
  • Noise has a significant impact on the performance of quantum computing. Quantum computers are inherently susceptible to errors and inaccuracies due to various sources of noise such as thermal environment and external control factors. The effects of noise affect different aspects of quantum computation. High gate error rates reduce the fidelity of quantum operations, affecting the accuracy of computations. Furthermore, the interaction with an environment causes decoherence, leading to the loss of quantum information. High decoherence levels limit the circuit depth that can be accurately executed on the quantum computer, directly affecting the complexity of the algorithms that can be executed. Furthermore, NISQ devices require periodic recalibration due to the effect of the thermal environment and the external control.
@@ III.2 A common language to discuss
The lack of a common and accepted standard for defining techniques and processes in quantum computing has led to the misuse of terminology. Even though there has been past attempts to classify the different types of benchmarks, and review approaches in the field, the terms quantum benchmarking, validation, verification, and testing have been loosely and inconsistently used in the literature. While we appreciate the groundwork laid by amico2023defining in establishing a foundation for benchmarking standards, our approach extends beyond insights solely derived from quantum computing research efforts. We aim to incorporate not only the lessons learned from the quantum domain but also draw upon the extensive findings from classical computing research. In this subsection, we will propose a unified terminology for the main concepts used in the quantum benchmarking field.
@@ III.3.1 Techniques
Mirror circuits: The concept of mirror circuits Proctor2021 ; PhysRevLett.129.150502 , inspired by Loschmidt echo, involves creating an alternative efficient classically simulatable circuit that mirrors a given quantum circuit $C$. This mirror circuit consists of preparing qubits in randomized Pauli eigenstates, running the base circuit $C$, applying a randomly selected Pauli layer $Q$, executing a reversed circuit $C^{-1}$ adjusted to implement a specific Pauli operation, and finally measuring qubits in their prepared basis. This approach is particularly effective for circuits containing Clifford and $Z(\theta)$ gates, a universal gate set widely used in quantum processors. The focus on Clifford gates –with $\theta$ being an integer multiple of $\pi/4$ in $Z(\theta)$ gates– is due to sufficiency in constructing various circuit types and the assumption of error-free implementation in the processors tested (verified processor), assuming error-free $Z(\theta)$ gates due to in-software phase updates. Through this method, each circuit in the set of mirror circuits has a straightforward and computable correct outcome, allowing the observed probabilities from these circuits to accurately represent the base circuit’s performance.
@@
Randomized Benchmarking (RB): Randomized Benchmarking (RB) is a technique used for characterizing the error rate of quantum processors. This technique involves applying a sequence of random gates, which can be sampled uniformly at random from the Haar distribution Emerson_2005 or uniformly at random from any unitary 2-design Dankert_2009 . RB has been employed as a technique for measuring errors, such as the average gate fidelity of Clifford gates PhysRevLett.106.180504 , the coherent or incoherent nature of the noise affecting Clifford gates Wallman_2015 , or the noise of a particular gate Magesan_2012 . While this technique is robust against state preparation and measurement noise, it only scales efficiently with the number of qubits when using Clifford gates.

### v3

@@ Abstract
The technological development of increasingly larger quantum processors on different quantum platforms raises the problem of how to fairly compare their performance, known as quantum benchmarking of quantum processors. This is a challenge that computer scientists have already faced when comparing classical processors, leading to the development of various mathematical tools to address it, but also to the identification of the limits of this problem. In this work, we briefly review the most important aspects of both classical processor benchmarks and the metrics comprising them, providing precise definitions and analyzing the quality attributes that they should exhibit. Subsequently, we analyze the intrinsic properties that characterize the paradigm of quantum computing and hinder the naive transfer of strategies from classical benchmarking. However, we can still leverage some of the lessons learned such as the quality attributes of a good benchmark such as relevance, reproducibility, fairness, verifiability and usability. Additionally, we review some of the most important metrics and benchmarks for quantum processors proposed in the literature, assessing what quality attributes they fulfill. Finally, we propose general guidelines for quantum benchmarking. These guidelines aim to pave the way for establishing a roadmap towards standardizing the performance evaluation of quantum devices, ultimately leading to the creation of an organization akin to the Standard Performance Evaluation Corporation (SPEC).
@@ I Introduction
Although quantum computers have not yet demonstrated a practical advantage over classical systems in real-world applications, this does not preclude the development of meaningful benchmarks. Moreover, there is theoretical evidence that quantum computers can offer significant advantages for certain problems Shor1994 . An analogous situation occurred during the early years of classical computing. During the early development of classical computing, benchmarks were primarily focused on evaluating low-level system, such as Whetstone WhetstoneBook and Dhrystone DhrystonePaper , rather than application-specific tasks. These benchmarks played a critical role in guiding hardware development and assessing system maturity.
In the early stages of classical computing, the lack of standardized benchmarking rules allowed manufacturers to define their own benchmark, often leading to both unintentional and strategic biases. In some cases, manufacturers even aggressively optimized their compilers or CPUs to perform well on specific benchmarks Hennessy2012 . This trend contributed to the emergence of a ”benchmarking industry” driven more by marketing considerations than by objective performance evaluation.
The field of quantum computing is similarly vulnerable. The lack of standardized benchmarking rules, together with the increasing competition among major industry players vying to outperform their rivals in the quantum computing race, might lead to a situation that resonates with Goodhart’s law goodhart1975problems : ”Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes”. In other words, the absence of a rigorously formulated and scrutinized approach to benchmarking in quantum computing may distort research priorities and hinder the development of truly scalable and reliable quantum processors. This bias occurs due to benchmarking models playing a crucial role on favouring what portion of the total ”behavior space” is captured SkadronChallengesInComputerArchitectureEvaluation .
@@ II What can we learn from Classical benchmarking?
A natural question that then arises is what we mean by performance. Performance is typically conceived in the classical computing scope as the useful work that a computer can get done in a given amount of time, and/or with a set of given resources. Indeed, in the past, it was common to hold the premise that ”the time required to perform a specified amount of computation is the ultimate measure of computer performance” CharacterizingComputerPerformanceWithASingleDigit . For classical computers, there have been historically two major approaches to the performance analysis problem: (1) Experimenting and measuring the actual system, and (2) Experimenting with a model of the system. For the latter, for any kind of performance evaluation not limited to computing devices you can use a physical model or a mathematical model. For computational performance assessment, we prefer to build models that can represent the system in terms of quantitative and logical relations. These models can later be either solved using mathematical tooling to obtain an analytical solution, or simulated Zeigler2018fg ; SimulationModelingAndAnalysisLawAverill . Even though in this work we will focus on the measurements of experiments with the actual systems, it is possible that we will see other approaches involving system modeling in the future. In fact, model-based approaches have been used by the quantum community for the assessment of quantum networks Wu2021 ; Coopmans2021 and the characterization of both incoherent and coherent errors in quantum gates Tripathi . Furthermore, other event-based models of quantum phenomena have also been published DeRaedt2005 ; Willsch2020 .
@@ II.1 Characteristics of a good benchmark
To be able to build a good benchmark, we introduce below a list of quality traits to ensure that it is comprehensive, accurate, and representative of system performance. These attributes are not intended as strict definitions, but rather allow for a degree of flexibility, making them applicable across a wide range of contexts. This approach aligns with the benchmarking principles proposed by Kounev Kounev2020mh and also reflect the most desirable characteristics of benchmarks proposed by other authors in the field HowToBuildABenchmark ; GustafsonAndSnell_HINT ; Henning_SpecCpu2000 ; Huppler_TheArtOfBuildingGoodBenchmark ; Sim_BenchmarkingToAdvanceResearch .
@@
Similarly, verifiability presents comparable challenges that we should carefully consider too. In industry, benchmarks are typically run by vendors who have a vested interest in the outcomes. In contrast, in academia results are subjected to a peer review process to be repeated and built upon by other researchers. Good benchmarks should introduce self-validation, since this may improve verifiability and usability, ensuring that results can be independently corroborated and effectively utilized.
@@
These statements highlight the susceptibility of benchmarks to manipulation, emphasizing the need for cautious interpretation and critical evaluation when drawing conclusions based on the results from ratio games. Other researchers such as Mashey have also call out the difficulties of aggregating data, and the controversy of using one Pythagorean mean over the others, as well as stated the common agreement that ”performance is not a single number” Mashey2004 . Nonetheless, even though there is agreement on the fact that a single number is delusional, there are still discussions to determine what is the preferred metric, sparking debates over which numerical representations are deemed more optimal in each context.
Finally, it is important to remember that while a good benchmark can provide a reliable and accurate measure of system performance, it cannot solve the problem of evaluating conflicting aspects of a computer device. Instead, it can help to properly balance the trade-offs for a particular objective and influence its strengths and weaknesses. It is not possible for a single benchmark to excel in all areas, and there is no silver bullet to solve this problem SkadronChallengesInComputerArchitectureEvaluation . Therefore, multiple workloads and benchmarks are necessary to comprehensively evaluate the performance of a system. Each benchmark may focus on different aspects of performance and may use different metrics and methodologies. At best, we can obtain a set of benchmarks that will be representative of the behavior space we want to model and measure. By combining the results of these multiple benchmarks, it would be possible to obtain a comprehensive and accurate picture of the system. Ultimately, the choice of benchmark(s) should depend on the specific use case and the objectives of the evaluation, as the relevance requirement already stated.
@@ II.2 What types of benchmarks are there?
Benchmarks can be classified according to multiple criteria, ranging from the objective of what portion of the behavior space they want to assess to the way the workload is defined. All of these categories fall under one of the following three different strategies when trying to assess the performance Lilja2005o ; Kounev2020mh :
@@ II.3 What makes a good metric
    Reliable. It must rank systems consistently with respect to a property that is subject to evaluation. In other words, if system $A$ performs better than system $B$ with respect to the property under evaluation, then the values of the metric for the two systems should consistently indicate so. The reliability of a metric is closely correlated with the relevance of a benchmark.
@@
    Consistent. It has the same units and the same precise definition across different systems or configurations of the same system. When this property is not met, the benchmark will suffer poor verifiability and potentially lack of relevance.
@@
Finally, it is important to highlight two key points. First, the properties discussed in this work are applicable not only to full-stack benchmarks and metrics but also to individual components of the computing stack. Second, the seemingly straightforward task of measuring a property using a metric might be tough. The effect of the Simpsons’ paradox serves as an adage to the designers of benchmarks and metrics, calling out the possibility of reversing or disappearing the trend on the data when combining it leboudec2010performance .
@@ II.4 Examples of classical benchmarks
Other examples of controversial benchmarks used historically are the synthetic ones, such as the Whetstone WhetstoneBook . Even though it was proposed to measure the floating point performance of a CPU, it can be misleading when compared with real-life situations. Nowadays, computers can easily fit the main loop of Whetstone in the low-level caches of the CPU, keeping the math coprocessor pipeline filled and achieving its optimal speed. Moreover, they face the issue of most synthetic benchmarks, they are not guaranteed to be correlated with real application scenarios, and fail to meet true relevance.
@@ III Quantum benchmarking
Quantum computers are devices that leverage the principles of quantum mechanics to address challenges beyond the computational capacity of classical computers. Several physical implementations for quantum computers have been explored, each employing different technologies as qubits, such as trapped ion QCion ; Zhang_2017 , superconducting circuits Arute2019 ; ZHU2022240 ; PhysRevLett.127.180501 , photonic QCPhoton ; PhysRevLett.127.180502 , and cold atoms Bernien2017ProbingMD ; Ebadi2020QuantumPO . Each of these architectures has its own set of advantages and disadvantages, and research is ongoing to determine which approach is most suitable for scaling up quantum computers to a practical level. Additionally, various quantum computing paradigms have emerged, including adiabatic quantum computing, which relies on adiabatic theorem to solve computational problems farhi2000 ; analog quantum computing, in which a controllable quantum device used to simulate other quantum systems Kendon_2010 ; digital quantum computing, which relies on the quantum circuit model for implementing general quantum algorithms doi:10.1126/science.273.5278.1073 ; and measurement-based quantum computing, where computation is performed through a sequence of measurements on a highly entangled state Raussendorf . These four paradigms offer different perspectives on the practical applications of quantum computing.
@@
In the quantum computing field, the term benchmarking has been used in a variety of contexts, sometimes leading to ambiguity. While it traditionally refers to a series of tests evaluating the performance of a device, its application has expanded to encompass the metrics evaluated as part of the test, the actual tests, or even for verifying the success of a test. We believe that establishing a unified and precise terminology for quantum benchmarking is essential to enable consistent and effective evaluation of quantum systems.
@@ III.1 Nuances of benchmarking quantum computers
    Fault Tolerant Quantum Computers (FTQC). Once quantum error correction codes can be efficiently implemented, it will be possible to run deeper and more complex quantum circuits without errors spreading uncontrollably. Although errors will not disappear completely, quantum information will be protected from the environment through correction schemes that limit the accumulation and spread of errors.
@@
    Fully Functional Quantum Computers (FFQC). Existing quantum algorithms and circuits consider the use of ancillary qubits, which play a similar role as cache memory in classical processors. However, for Fully Functional Quantum Computers, there will also be necessary the existence of Quantum Random Access Memories (qRAM) QuantumRam , and even long-term (error-corrected) quantum memories Terhal2015 .
@@
The challenges that we are facing today within the NISQ era are different from the ones we will be facing in the upcoming eras, and hence the benchmarks should reflect this. Given the wide diversity of technological platforms and paradigms for implementing quantum computers, it is likely that benchmarks in the NISQ era will focus on comparing the performance among different quantum computers within subsets of the broad existing variety. As quantum computing technology advances, it is likely that certain technological platforms will become dominant, particularly those that demonstrate scalability, fault tolerance, or economic viability for large-scale implementation. However, this convergence does not avoid the possibility of continued diversification. As seen in classical computing with the emergence of application-specific integrated circuits (ASICs), technological maturity can also drive specialization tailored to distinct tasks. Similarly, in quantum computing, different platforms may evolve to address specific application domains. In Sec. IV, we will discuss further what recommendations we suggest to properly face the different quantum technological development eras.
@@
  • The fundamental unit of information in a quantum computer is the qubit (quantum bit). While classical bits can only represent a single binary value 0 or 1, a qubit is described as a linear combination of the $|0\rangle$ and $|1\rangle$ states.
@@
Beyond these fundamental features associated to quantum physics, there are other relevant aspects of quantum computing which mainly affect the performance for the NISQ and PQEC eras.
@@
  • While errors in classical computing are extensively studied and managed through classical error correction protocols, noise significantly affects the performance of quantum computing systems. Quantum computers are inherently susceptible to errors and inaccuracies due to various sources of noise such as thermal environment and external control factors. The effects of noise affect different aspects of quantum computation. High gate error rates reduce the fidelity of quantum operations, affecting the accuracy of computations. Furthermore, the interaction with an environment causes decoherence, leading to the loss of quantum information. High decoherence levels limit the circuit depth that can be accurately executed on the quantum computer, directly affecting the complexity of the algorithms that can be executed. Furthermore, NISQ devices require periodic recalibration due to the effect of the thermal environment and the external control.
@@ III.2 A common language to discuss
The lack of a common and accepted standard for defining techniques and processes in quantum computing has led to the misuse of terminology. Even though there has been past attempts to classify the different types of benchmarks, and review approaches in the field, the terms quantum benchmarking, validation, verification, and testing have been loosely and inconsistently used in the literature. While we appreciate the groundwork laid by amico2023defining ; proctor2024benchmarkingquantumcomputers in establishing a foundation for benchmarking standards and properties, our approach builds upon not only insights from quantum computing research but also draws from the well-established and mature field of classical computing benchmarking. In addition, we support our proposal with numerous carefully analyzed examples, which help to illustrate and validate the concepts and classifications we introduce. In this subsection, we will propose a unified terminology for the main concepts used in the quantum benchmarking field.
@@ III.3.1 Techniques
Mirror circuits: The concept of mirror circuits Proctor2021 ; PhysRevLett.129.150502 , inspired by Loschmidt echo, involves creating an alternative efficient classically simulatable circuit that mirrors a given quantum circuit $C$. The mirror circuit technique generally consists of implementing a circuit $C$ followed by its inverse $C^{-1}$. There are several variations or specifications of this method. One common approach involves preparing the qubits in randomized Pauli eigenstates, executing the base circuit $C$, applying a randomly selected Pauli layer $Q$, running the reversed circuit $C^{-1}$ adjusted to implement a specific Pauli operation, and finally measuring the qubits in their initial basis. Another variation lies in how $C^{-1}$ is implemented, either by reversing all the gates in $C$, or by applying a different random transpilation of the circuit.
This approach is particularly effective for circuits containing Clifford and $Z(\theta)$ gates, with $\theta$ being an integer multiple of $\pi/4$ in $Z(\theta)$ gates, a universal gate set widely used in quantum processors. Through this method, each circuit in the set of mirror circuits has a straightforward and computable correct outcome, allowing the observed probabilities from these circuits to accurately represent the base circuit’s performance.
@@
Randomized Benchmarking (RB): Randomized Benchmarking (RB) is a technique used for characterizing the error rate of quantum processors. This technique involves applying a sequence of random gates, which can be sampled uniformly at random from the Haar distribution Emerson_2005 or uniformly at random from any unitary 2-design Dankert_2009 , such as the Clifford group. RB has been employed as a technique for measuring errors, such as the gate fidelity of Clifford gates PhysRevLett.106.180504 , the coherent or incoherent nature of the noise affecting Clifford gates Wallman_2015 , or the noise of a particular gate Magesan_2012 , among others. This technique is robust against state preparation and measurement (SPAM) noise and is considered scalable. Although the standard randomized benchmarking (RB) protocol employing this technique is theoretically scalable, in practice it becomes challenging to implement for more than 5 or 6 qubits due to the depth of the required Clifford circuits. To overcome this limitation, recent variants of the protocol have been developed that aim to minimize circuit depth, making the method viable for systems with up to 20 qubits Hines2024 ; PhysRevLett.129.150502 ; polloreno2023 . For a detailed review of the different types of protocols based on this technique, see hashim2024 .